OpenStudy (anonymous):

f(x)=1/4(x+1)^2+6 find vetex,line of symmetry,max/min of f(x), value of f(-1)=6 the max or min? having a lot of trouble with this because of the 1/4

OpenStudy (rulnick):

vertex is at x= -1, y=6

OpenStudy (rulnick):

line of symm is x=-1

OpenStudy (rulnick):

min is at vertex (f(x)=6)

OpenStudy (rulnick):

value at f(-1)=6 is min

OpenStudy (rulnick):

... get all this from a=1/4, b=1/2, c=25/4, after expanding the (x+1)^2 and simplifying.

OpenStudy (anonymous):

so 6 is the max?

OpenStudy (rulnick):

No, a min. The min value of the function is 6, meaning f(x)=6, and it occurs at x=-1.